OpenStudy (anonymous):

will fan and medal. will fan and medal will fan and medal. will fan and medal !!!!!!!!!!!!!!!!!!!!!!question down blow!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

whats question?

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

which of the following expressions is equal to \[\sqrt[7]{k}\] \[7^{1/k}\] \[(\frac{ 1 }{ 7 })^{k}\] \[(\frac{ 1 }{ k })^{7}\] \[k ^{\frac{ 1 }{ 7 }}\]

OpenStudy (anonymous):

@rebeccaxhawaii @123ABC456C

OpenStudy (anonymous):

@123AB456C

OpenStudy (anonymous):

will this help? https://www.math.byu.edu/~klkuttle/LinearAlgebraSolutions.pdf

rebeccaxhawaii (rebeccaxhawaii):

\[\sqrt[a]{x}=x ^{1/a}\]

OpenStudy (anonymous):

yasssssssssss I forgot that thanks

rebeccaxhawaii (rebeccaxhawaii):

no problem so which one would it be?